can you give an example of how to think on any advance level physics problems conceptually. because i can understand the physics but how to apply it to the problems can you quote an example on circular motionn

Grade:

1 Answers

AJIT AskiitiansExpert-IITD
13 years ago

Dear  seema ,

First of all make a diagram of the question and label it. Then , see what is asked and try to remember the related equations to it and write them . eliminate the unknowns and find the answer. Actually it is just this much easy if you remain careful at every step of it , no need to confuse yourself by adopting techniques . Just be confident and apply the simple procedure. 

as an examplae suppose you are asked to find the tension at the highest point when a body is moved in a vertical circle ,

then make a diagram and label it . then see when tension is asked you have to use force equation at the highest point .

T + mg  =  mv2 / r

now you dont know V here .

For that apply energy conservaiton ,

2mgr = 1/2( mu2 -mv2 ) and find the value of v at the highest point given the initial velocity u . put there and get the answer.

 

HOPE it helps!!!!

 

 

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

Askiitians Expert

Ajit Singh Verma IITD